OpenStudy (anonymous):

The current I in an electrical conductor varies inversley as the resistance R of the conductor. The current is 3 amperes when the resistance is 990ohms. What is the current when the resistance is 540ohms

OpenStudy (anonymous):

for inverse relations xy = constant here IR = V (here V is the constant) I1R1 = 3 * 990 = 2970 So I2R2 also = 2970 I2* 540 = 2970 I2 = 2970/540 I2 = 5.5 so current in second case is 5.5 A

OpenStudy (anonymous):

i = k/R 3 = k / 990 k = 2970 so i = 2970 / R when R = 540 current i = 2970 / 540 = 5.5 amperes